### v1

# On the asymptotic behavior of the spectral gap for discrete Schrödinger operators

Abstract.
In this note we elaborate on the convergence of the spectral gap of a class of discrete Schrödinger operators defined on a path graph in the limit of infinite volume. We confirm and extend recent results on the convergence of the spectral gap to a larger class of potentials using entirely different methods. Most importantly, we also resolve a conjecture previously proposed in this context. This then yields new insights into the rate at which the spectral gap tends to zero as the volume increases.
2010 Mathematics Subject Classification:
1. Introduction
This paper is devoted to the investigation of the spectral gap of discrete Schrödinger operators. Here, spectral gap refers to the difference between the lowest two eigenvalues. At least in the continuous setting, related considerations have a long tradition: for example, important results were obtained in [AB89] and then in [Lav94], solving the one-dimensional version of the fundamental gap conjecture [AC11]. As a matter of fact, it turns out that it is in general quite difficult to say something about how the gap changes under perturbations, for instance, under the addition of some non-negative potential. As a matter of fact, it might already be difficult to infer the correct sign of the changes [Abr91]. This is certainly true as long as the configuration space on which the operator is defined remains fixed; for example, one could think of a Schrödinger operator defined on an interval of fixed length [AK20, ACH21]. However, as demonstrated recently [KT23, KT24], one is able to say more about the spectral gap for a larger class of potentials if one is working with Schrödinger operators defined on domains of increasing volume. Most importantly, in [KT24] the authors discussed an interesting effect which was recently studied on the path graph [KY25] and which forms – in some sense – the starting point of our investigations. More explicitly, it was shown that the presence of (certain) compactly supported potentials drastically change the asymptotic behavior of the spectral gap when compared to the asymptotic behavior of the gap of the Laplacian without a potential. In other words, as soon as one adds such a non-negative compactly supported potential to the Laplacian, the spectral gap converges strictly faster to zero in the limit of infinite volume. This somewhat surprising effect is the result of an effective degeneracy of the lowest two eigenvalues at infinite volume. It is important to note, however, that this effect might disappear in higher dimensions or reappear as well, depending on the particular nature of the potential [KT23]. This also explains why the authors of [KY25] focused on a specific discrete graph – the path graph: adding a compactly supported potential to the Laplacian then leads, at least on an intuitive level, to the mentioned (approximate) degeneracy of the lowest two eigenvalues in the infinite-volume limit since the potential cuts the graph into two (approximately) congruent pieces.
Starting with the results of [KY25], we want to achieve two things with this paper: First, since the authors of [KY25] considered only a very specific potential localized on only one vertex, we generalize their main result [KY25, Theorem 6] to general non-negative and compactly supported potentials, while using entirely different proof techniques. Furthermore, we prove the conjecture put forward in [KY25, Section 4] which provides more information on how fast the spectral gap converges to zero. By doing this, we provide some rigorous justification to the numerical results obtained in [KY25, Section 4].
For completeness, let us also mention that investigations of the spectral gap (or spectral properties) for Jacobi operators [Moe76, Tes00] and Schrödinger-type operators on graph-like structures such as combinatorial graphs [KLW21] and metric graphs [BK13, Kur24] are quite common. Regarding the discrete case, let us refer to [FCLP18, LS20, JMM22, BS24]; for the metric case, one may consult [KGMN13, KKMM16, BL17, Roh17, BKKM17] and references therein.
Our paper is organized as follows: In Section 2 we introduce the basic setting and the class of Schrödinger operators considered. In Section 3 we then derive upper and lower bounds on the two lowest eigenvalues. Finally, in Section 4 we use these estimates to prove our main result (Theorem 7) and to resolve a conjecture from [KY25, Section 4] (Corollary 9).
2. The model
Our configuration space is the path graph $G_{k}=(V_{k},E_{k})$ with vertex set $V_{k}:=\{-k,-k+1,\ldots,k\}$ and hence with an odd number of vertices $|V_{k}|=2k+1$, $k\in\mathds{N}=\{1,2,$...\}. Except for the outer two vertices that have one neighbor each, it is readily clear that all other vertices have exactly two neighbors. This implies that the edge set is given by $E_{k}=\{\{v,w\}:v,w\in V_{k}\ \text{with}\ |v-w|=1\}$.
In the following, we study operators defined on the finite-dimensional Hilbert space $\mathcal{H}_{k}=\mathds{C}^{|V_{k}|}$, $k\in\mathds{N}$. The discrete (standard, unweighted) Laplacian $\mathcal{L}_{\gamma,k}:\mathds{C}^{|V_{k}|}\rightarrow\mathds{C}^{|V_{k}|}$ is defined via
$(\mathcal{L}_{k}f)(v):=\sum_{w\in V_{k}}\gamma_{w,v}(f(v)-f(w))\ ,\quad v\in V_{k}\ ,\quad f\in\mathcal{H}_{k},\quad k\in\mathds{N}\ ,$
where $\gamma_{v,w}=1$ for all $v,w\in V_{k}$ with |vw|=1|v-w|=1 and $\gamma_{v,w}=0$ else. The associated quadratic form is given by
$q_{k}[f]:=\frac{1}{2}\sum_{v,w\in V_{k}}\gamma_{w,v}|f(v)-f(w)|^{2}=\frac{1}{2}\sum_{v,w\in V_{k}}\gamma_{w,v}|\nabla_{v,w}f|^{2}\ ,\quad f\in\mathcal{H}_{k}\ ,$
where we defined $\nabla_{v,w}f:=f(v)-f(w)$. Since the Laplacian $\mathcal{L}_{k}$ is a self-adjoint (non-negative) operator on a finite-dimensional Hilbert space, its spectrum consists of (non-negative, real) eigenvalues only. Furthermore, inserting the vector $\mathds{1}=(1,...,1)^{T}\in\mathcal{H}_{k}$ into $q_{k}[\cdot]$ shows that the lowest eigenvalue of $\mathcal{L}_{k}$, for all $k\in\mathds{N}$, is zero.
In a next step we construct a discrete Schrödinger operator by introducing a non-negative external potential with compact support. More explicitly, the quadratic form of our Schrödinger operator shall be given by
$q_{k,\boldsymbol{\alpha}}[f]:=\frac{1}{2}\sum_{v,w\in V_{k}}\gamma_{w,v}|\nabla_{v,w}f|^{2}+\sum\limits_{j\in J}\alpha_{j}|f(j)|^{2}\ ,\quad f\in\mathcal{H}_{k}\ ,$ (1)
where $J\subset\mathds{Z}$ is a non-empty set independent of k; $\alpha_{j}>0$, $j\in J$, represents the strength of the external potential localized at the $j$-th vertex. The associated self-adjoint operator then becomes
$H_{k,\boldsymbol{\alpha}}:=\mathcal{L}_{k}+\sum\limits_{j\in J}\alpha_{j}\delta_{j}\ ,\quad k\in\mathds{N}\ ,$ (2)
where $\boldsymbol{\alpha}=(\alpha_{j})_{j\in J}$ and $\delta_{j}$ is such that $(\delta_{j}f)_{i}:=\delta_{ij}f(j)$, $i\in V_{k}$, $\delta_{ij}$ referring to the Kronecker delta. As in the case for $\mathcal{L}_{k}$, the operator $H_{k,\boldsymbol{\alpha}}$ is, for all $k\in\mathds{N}$, non-negative and its spectrum consists of non-negative eigenvalues only. For all $k\in\mathds{N}$, we denote the eigenvalues of $H_{k,\boldsymbol{\alpha}}$ by
$\lambda_{0}(V_{k},\boldsymbol{\alpha})<\lambda_{1}(V_{k},\boldsymbol{\alpha})\leq...\leq\lambda_{|V_{k}|-1}(V_{k},\boldsymbol{\alpha})\ .$
We can now introduce the central object of interest in this paper, which is the spectral gap
$\Gamma(V_{k},\boldsymbol{\alpha}):=\lambda_{1}(V_{k},\boldsymbol{\alpha})-\lambda_{0}(V_{k},\boldsymbol{\alpha})\ .$ (3)
Note that $\Gamma(V_{k},\boldsymbol{\alpha})>0$ for each value of $k\in\mathds{N}$, since the ground state is non-degenerate; more on the spectral theory of graphs can be found in [Chu97, BH12, KLW21]. For convenience, we set
$\widetilde{\alpha}\coloneq\min\limits_{j\in J}\alpha_{j}\qquad\text{ and }\qquad\widehat{\alpha}\coloneq\sum\limits_{j\in J}\alpha_{j}\ .$
It is important to mention that for the special case $J=\{0\}$ with external potential $\alpha\delta_{0}$, one has
$\lambda_{1}(V_{k},\alpha\delta_{0})=\lambda_{0}(V_{k},\infty\delta_{0})$ (4)
for all $\alpha\geq 0$. Here, $\lambda_{1}(V_{k},\alpha\delta_{0})$ is shorthand for $\lambda_{1}(V_{k},\boldsymbol{\alpha})$ with external potential $\alpha\delta_{0}$ and $\lambda_{0}(V_{k},\infty\delta_{0})$ is shorthand for the lowest eigenvalue of the Laplacian on the path graph with a Dirichlet condition at the zero vertex. Later we shall also use that there exist constants $c_{1},c_{2}>0$ such that for all but finitely many $k\in\mathds{N}$,
$c_{1}k^{-2}\leq\lambda_{0}(V_{k},\infty\delta_{0})\leq c_{2}k^{-2}\ ;$ (5)
this follows from a well-known explicit formula for $\lambda_{1}(V_{k},\alpha\delta_{0})$ as used, for example, in [KY25]. In the following, we will also frequently use the identity
$\frac{1}{2}\sum_{v,w\in V_{k}}\gamma_{w,v}|\nabla_{v,w}f|^{2}=\sum_{j=-k}^{k-1}|\nabla_{j}f|^{2}=\sum_{j=-k}^{k}|\nabla_{j}f|^{2}$
where $\nabla_{j}f:=f(j+1)-f(j)$ and where we set $f(k+1):=f(k)$.
We conclude this section with a remark to help the reader better understand our notation used in the following.
Remark 1.
Firstly, throughout this paper, we denote by $\varphi_{k}$ the normalized ground state of $H_{k,\boldsymbol{\alpha}}$, that is, the eigenvector corresponding to the lowest eigenvalue $\lambda_{0}(V_{k},\boldsymbol{\alpha})$ with $\|\varphi_{k}\|^{2}_{\mathcal{H}_{k}}=\sum_{j=-k}^{k}|\varphi_{k}(j)|^{2}=1$. Note that the ground state $\varphi_{k}$ can be chosen to be strictly positive.
Secondly, for two sequences $(a_{k})_{k\in\mathds{N}}$, $(b_{k})_{k\in\mathds{N}}$, $a_{k}\neq 0$ for all but finitely many $k\in\mathds{N}$, we use the notation $b_{k}\lesssim a_{k}$ iff there is a constant $c>0$ such that $|b_{k}/a_{k}|\leq c$ for all but finitely many $k\in\mathds{N}$, and $b_{k}\sim a_{k}$ iff there are constants $c_{1},c_{2}>0$ such that $c_{1}\leq|b_{k}/a_{k}|\leq c_{2}$ for all but finitely many $k\in\mathds{N}$.
Also, in the proofs we denote by $(const.)$ positive constants that are independent of $k$ and the potential, whenever their precise values are irrelevant to the core arguments of our proofs.
3. Spectral estimates
In this section we provide upper and lower bounds for $\lambda_{0}(V_{k},\boldsymbol{\alpha})$ as well as for $\lambda_{1}(V_{k},\boldsymbol{\alpha})$. We start with a lower bound for $\lambda_{0}(V_{k},\boldsymbol{\alpha})$. We set
$r_{\min}\coloneq\min\{j\in J\}\qquad$ $\text{ and }\qquad r_{\max}\coloneq\max\{j\in J\}\ ,$
and introduce
$a_{k,1}\coloneq 1/2-\sum_{j=-k}^{r_{\min}}|\varphi_{k}(j)-\varphi_{k}(r_{\min})|^{2}\quad$ $\text{ and }\quad a_{k,2}\coloneq 1/2-\sum_{j=r_{\max}}^{k}|\varphi_{k}(j)-\varphi_{k}(r_{\max})|^{2}\ .$
Note that $1\geq a_{k,1}+a_{k,2}\geq 1-\sum_{j=-k}^{k}|\varphi_{k}(j)|^{2}=0$ due to monotonicity of the ground state outside of $r_{\min},r_{\max}]$.
Theorem 2 (Lower bound ground state energy).
Consider a Hamiltonian $H_{k,\boldsymbol{\alpha}}$ with ground state energy $\lambda_{0}(V_{k},\boldsymbol{\alpha})$. Then, for all $k>\max\{|r_{\min}|,|r_{\max}|\}/2$, we have
$\lambda_{0}(V_{k},\boldsymbol{\alpha})$ $\geq\Big{(}\dfrac{1}{2}-a_{k,1}\Big{)}\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0})+\Big{(}\dfrac{1}{2}-a_{k,2}\Big{)}\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})\ .$ (6)
Moreover,
$0\leq a_{k,1}+a_{k,2}\lesssim\widetilde{\alpha}^{-1}k^{-1}\ .$ (7)
Proof.
In a first step we obtain
$\lambda_{0}(V_{k},\boldsymbol{\alpha})$ $\geq\sum\limits_{j=-k}^{r_{\min}-1}|\nabla_{j}(\varphi_{k}-\varphi_{k}(r_{\min}))|^{2}+\sum\limits_{j=r_{\max}}^{k}|\nabla_{j}(\varphi_{k}-\varphi_{k}(r_{\max}))|^{2}+\sum\limits_{j\in J}\alpha(j)|\varphi_{k}(j)|^{2}$
$\geq\sum\limits_{j=-k}^{r_{\min}}|\varphi_{k}(j)-\varphi_{k}(r_{\min})|^{2}\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0})$
$\qquad+\sum\limits_{j=r_{\max}}^{k}|\varphi_{k}(j)-\varphi_{k}(r_{\max})|^{2}\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})+\sum\limits_{j\in J}\alpha(j)|\varphi_{k}(j)|^{2}\ .$
Using the definition of $a_{k,1},a_{k,2}$ and the fact that the last term is non-negative then implies the lower bound (6).
We now proceed to prove (7): Since $2\lambda_{0}(V_{k},\boldsymbol{\alpha})\leq\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0})+\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})$, we conclude
$\sum\limits_{j\in J}\alpha(j)|\varphi_{k}(j)|^{2}\leq a_{k,1}\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0})+a_{k,2}\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})\ ,$ (8)
and
$\sum\limits_{j\in J}|\varphi_{k}(j)|^{2}\leq\widetilde{\alpha}^{-1}\Big{(}a_{k,1}\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0})+a_{k,2}\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})\Big{)}\ .$ (9)
Next, we observe that
$1=\sum\limits_{j=-k}^{k}|\varphi_{k}(j)|^{2}=\sum\limits_{j=-k}^{r_{\min}-1}|\varphi_{k}(j)|^{2}+\sum\limits_{j=r_{\max}+1}^{k}|\varphi_{k}(j)|^{2}+\sum\limits_{j=r_{\min}}^{r_{\max}}|\varphi_{k}(j)|^{2}$
and therefore
$\begin{split}a_{k,1}+a_{k,2}&=\sum\limits_{j=r_{\min}}^{r_{\max}}|\varphi_{k}(j)|^{2}+2\varphi_{k}(r_{\min})\sum\limits_{j=-k}^{r_{\min}-1}\varphi_{k}(j)+2\varphi_{k}(r_{\max})\sum\limits_{j=r_{\max}+1}^{k}\varphi_{k}(j)\\ &\qquad\qquad-\varphi_{k}^{2}(r_{\min})(k+r_{\min})-\varphi_{k}^{2}(r_{\max})(k-r_{\max})\ .\end{split}$ (10)
To simplify we set $\Theta_{\max}:=\max\big{\{}\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0}),\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})\big{\}}$. Then, using (9) and Hölder’s inequality as well as $a+b\leq\max\{2a,2b\}$ for $a,b>0$, we get
$a_{k,1}+a_{k,2}$ $\leq(const.)\max\Big{\{}\widetilde{\alpha}^{-1/2}\left(a_{k,1}+a_{k,2}\right)^{1/2}\Theta_{\max}^{1/2}k^{1/2},\sum\limits_{j=r_{\min}}^{r_{\max}}|\varphi_{k}(j)|^{2}\Big{\}}\ .$
In order to estimate the last sum, we use that for all $j\in\{r_{\min},\ldots,r_{\max}\}$, again due to (9) and the eigenvalue equation,
$|\varphi_{k}(j)|^{2}$ $\leq\left(|\varphi_{k}(r_{\min})|+\sum_{i=r_{\min}}^{j-1}|\varphi_{k}(i+1)-\varphi_{k}(i)|\right)^{2}$ (11)
$\leq 2\max_{j\in J}|\varphi_{k}(j)|^{2}+2|r_{\max}-r_{\min}|\lambda_{0}(V_{k},\alpha)\ .$
Therefore, using $a+b\leq\max\{2a,2b\}$ for $a,b>0$, we get
$\begin{split}a_{k,1}+a_{k,2}&\leq(const.)\max\Big{\{}\widetilde{\alpha}^{-1}\Theta_{\max}k,\Theta_{\max}\Big{\}}\ .\end{split}$ (12)
The claimed estimate (7) now follows by observing that $\Theta_{\max}\sim k^{-2}$.
Remark 3.
Inspecting the proof of Theorem 2, in particular (8), we obtain the following estimate for the special case when $J=\{0\}$, that is, for $H_{k,\alpha}=\mathcal{L}_{k}+\alpha\delta_{0}$ with $\alpha>0$: There exists a constant $c>0$ such that, for all but finitely many $k\in\mathds{N}$,
$\varphi_{k}(0)\leq c\alpha^{-1}k^{-3/2}\ .$
This can be regarded as a generalization of [KY25, Lemma 5]. Note here that $\varphi_{k}(0)=(2k+1)^{-1/2}$ for $\alpha=0$.
From a physics point of view, it is interesting to remark that the just mentioned estimate implies an estimate on the potential energy of a particle described by $H_{k,\alpha}=\mathcal{L}_{k}+\alpha\delta_{0}$ with $\alpha>0$: More explicitly, the potential energy in the ground state is given by $E_{pot}(\alpha):=\alpha|\varphi_{k}(0)|^{2}$ and we therefore obtain, for $k\in\mathds{N}$,
$E_{pot}(\alpha)\leq c^{2}\alpha^{-1}k^{-3}$
with the constant $c>0$ from above. The interesting point here is that the potential energy is of a lower order than the kinetic energy, which is of order $\sim k^{-2}$.
Remark 4.
We encourage the reader to keep the following physical intuition in mind, which also served as a guiding perspective in the development of this work: Our focus lies on the two lowest eigenvalues, both of which converge to zero as the size of the graph tends to infinity. In this regime, a particle occupying a quantum state associated with either of these eigenvalues has arbitrarily low energy. As a result, any potential of finite strength effectively appears to be of an infinite height to such a low-energy particle. As a consequence, this provides some explanation for the emergence of quantities in our statements related to infinitely strong potentials.
In a next statement we prove an upper bound for the lowest eigenvalue $\lambda_{0}(V_{k},\boldsymbol{\alpha})$. In order to formulate it, we introduce the two functions
$\begin{split}\widetilde{\cos}_{k,r_{\min}}(j)&\coloneq\begin{cases}A_{\min}^{-1/2}\cos\left(\left[j+k+\frac{1}{2}\right]\pi/(2(k+r_{\min})+1)\right),\ -k\leq j\leq r_{\min}\ ,\\ 0\ \quad\text{else}\ ,\end{cases}\\ \widetilde{\cos}_{k,r_{\max}}(j)&\coloneq\begin{cases}A_{\max}^{-1/2}\cos\left(\left[k-j+\frac{1}{2}\right]\pi/(2(k-r_{\max})+1)\right),\ r_{\max}\leq j\leq k\ ,\\ 0\ \quad\text{else}\ .\end{cases}\end{split}$
Here, $A_{\max},A_{\min}>0$ are such that both functions are normalized to $1/2$. Also note that both functions are constructed using the well-known explicit form of the second eigenfunction to the Hamiltonian $H_{k,\boldsymbol{\alpha}}=\mathcal{L}_{k}+\alpha\delta_{0}$.
Theorem 5 (Upper bound ground state energy).
Consider a Hamiltonian $H_{k,\boldsymbol{\alpha}}$ with ground state $\varphi_{k}\in\mathcal{H}_{k}$ corresponding to the ground state energy $\lambda_{0}(V_{k},\boldsymbol{\alpha})$. Then
$\lambda_{0}(V_{k},\boldsymbol{\alpha})\leq\dfrac{1}{2}(1-b_{k})\Big{(}\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0})+\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})\Big{)}+b_{k}c_{k}^{2}\ ,$ (13)
where $c_{k}^{2}:=\lambda_{0}(V_{k},\infty\delta_{0})/(2+\epsilon)$ for an arbitrary $\epsilon>0$ and with the sequence $(b_{k})_{k\in\mathds{N}}$, $0<b_{k}\leq 1$ for all $k\in\mathds{N}$, such that
$\psi_{k}(j):=\left(1-b_{k}\right)^{1/2}\left(\widetilde{\cos}_{k,r_{\min}}(j)+\widetilde{\cos}_{k,r_{\max}}(j)\right)+\widehat{\alpha}^{-1/2}b^{1/2}_{k}c_{k}\ ,\quad j\in V_{k}\ ,\quad k\in\mathds{N}\ ,$
is normalized. Furthermore, there exists a constant $c_{1}>0$ independent of $k$ and $\boldsymbol{\alpha}$ such that
$b_{k}\geq c_{1}\widehat{\alpha}^{-1}k^{-1}$ (14)
for all but finitely many $k\in\mathds{N}$.
Proof.
We set $\widehat{c}_{k}\coloneq\widehat{\alpha}^{-1/2}c_{k}$ and write
$\psi_{k}(j)=\left(1-b_{k}\right)^{1/2}\left(\widetilde{\cos}_{k,r_{\min}}(j)+\widetilde{\cos}_{k,r_{\max}}(j)\right)+b^{1/2}_{k}\widehat{c}_{k}\ ,\quad j\in V_{k}\ ,\quad k\in\mathds{N}\ .$
Calculating the norm of $\psi_{k}$ gives the relation
$b_{k}$ $=2\left(1-b_{k}\right)^{1/2}b_{k}^{1/2}\widehat{c}_{k}\sum\limits_{j=-k}^{k}\left(\widetilde{\cos}_{k,r_{\min}}(j)+\widetilde{\cos}_{k,r_{\max}}(j)\right)+(2k+1)b_{k}\widehat{c}_{k}^{2}\ .$
This equation can be solved explicitly for $b_{k}$, yielding the desired sequence. Recall that $(b_{k})_{k\in\mathds{N}}$ is such that $0<b_{k}\leq 1$ for all $k\in\mathds{N}$ and $\|\psi_{k}\|_{\mathcal{H}_{k}}=1$ for all $k\in\mathds{N}$.
Employing the minmax-principle using $\psi_{k}$ as a trial state then implies
$\lambda_{0}(V_{k},\boldsymbol{\alpha})$ $\leq(1-b_{k})\dfrac{1}{2}\Big{(}\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0})+\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})\Big{)}+b_{k}c_{k}^{2}\ .$
Comparing the upper bound for $\lambda_{0}(V_{k},\boldsymbol{\alpha})$ with the lower bound in Theorem 2, we conclude that, for all but finitely many $k\in\mathds{N}$, $b_{k}\leq(const.)(a_{k,1}+a_{k_{2}})$ and therefore $\lim_{k\to\infty}b_{k}=0$ by Theorem 2.
Consequently, for all but finitely many $k\in\mathds{N}$,
$b_{k}$ $\geq b_{k}^{1/2}\widehat{c}_{k}\sum\limits_{j=-k}^{k}\big{(}\widetilde{\cos}_{k,r_{\min}}(j)+\widetilde{\cos}_{k,r_{\max}}(j)\big{)}\ .$
Hence, using that $\sum\limits_{j=-k}^{k}\big{(}\widetilde{\cos}_{k,r_{\min}}(j)+\widetilde{\cos}_{k,r_{\max}}(j)\big{)}\geq(const.)k^{1/2}$ for all but finitely many $k\in\mathds{N}$ and (5), we obtain
$b_{k}\geq(const.)\widehat{\alpha}^{-1}k^{-1}\ .$
We conclude this section by providing upper and lower bounds for the second-lowest eigenvalue $\lambda_{1}(V_{k},\boldsymbol{\alpha})$.
Lemma 6 (Bounds on the first excited state energy).
For all but finitely many $k\in\mathds{N}$ one has
$\lambda_{0}(V_{k},\infty\delta_{0})\leq\lambda_{1}(V_{k},\boldsymbol{\alpha})\leq\max\big{\{}\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0}),\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})\big{\}}\ .$
Proof.
The lower bound is a direct consequence of $\lambda_{1}(V_{k},\boldsymbol{\alpha})\geq\lambda_{1}(V_{k},0)=\lambda_{0}(V_{k},\infty\delta_{0})$. The upper bound follows from the operator inequality $\mathcal{L}_{k}+\sum_{j\in J}\alpha_{j}\delta_{j}\leq\mathcal{L}_{k}+\sum_{j\in J}\infty\delta_{j}$ in combination with $\lambda_{1}(V_{k},\sum_{j\in J}\infty\delta_{j})\leq\max\{\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0}),\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})\}$ that holds for all but finitely many $k\in\mathds{N}$. ∎
4. Main results
In this section, we study properties of the sequence
$\left(|V_{k}|^{3}\cdot\Gamma(V_{k},\boldsymbol{\alpha})\right)_{k\in\mathds{N}}$
and prove a conjecture regarding its limiting behavior as put forward in [KY25]. By doing this and by considering potentials of compact support, we also generalize the main result of [KY25] – both – in terms of convergence speed and applicability to a broader class of potentials: more explicitly, we will show that
$\lim_{k\rightarrow\infty}\left(|V_{k}|^{2+\eta}\cdot\Gamma(V_{k},\boldsymbol{\alpha})\right)=0$
for all $\eta<1$, as soon as the external potential is not the zero potential. Here, one should recall that $\lim_{k\rightarrow\infty}\left(|V_{k}|^{2}\cdot\Gamma(V_{k},\boldsymbol{\alpha})\right)=\pi^{2}$ for $\boldsymbol{\alpha}=0$; compare with [KY25]. In other words, as soon as there is a non-vanishing external potential of compact support, the spectral gap converges strictly faster to zero than in the absence of any potential. This might be surprising at first, since the potential is supported on a smaller and smaller fraction of the configuration space in the limit of infinite volume.
We now state the main result of this paper.
Theorem 7 (Asymptotic upper and lower bounds I).
Consider a Hamiltonian $H_{k,\boldsymbol{\alpha}}=\mathcal{L}_{k}+\sum_{j\in J}\alpha_{j}\delta_{j}$ with compactly supported external potential and associated spectral gap $\Gamma(V_{k},\boldsymbol{\alpha})$. Then, the following holds:
  1. (i)
    There exists a constant $c_{1}=c_{1}(\boldsymbol{\alpha})>0$ such that
    $|V_{k}|^{3}\cdot\Gamma(V_{k},\boldsymbol{\alpha})\leq c_{1}\quad\text{for all}\quad k\in\mathds{N}\ .$
  2. (ii)
    If $\widehat{\alpha}=\sum_{j\in J}\alpha_{j}$ is sufficiently small or if $J=\{0\}$, there exists a constant $c_{2}=c_{2}(\boldsymbol{\alpha})>0$ such that
    $|V_{k}|^{3}\cdot\Gamma(V_{k},\boldsymbol{\alpha})\geq c_{2}\quad\text{for all}\quad k\in\mathds{N}\ .$
Proof.
To simplify notation, we set $\Theta_{\max}:=\max\big{\{}\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0}),\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})\big{\}}$ and $\Theta_{\min}:=\min\big{\{}\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0}),\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})\big{\}}$.
We start with (i): Theorem 2 and Lemma 6 immediately imply for all but finitely many $k\in\mathds{N}$,
$|V_{k}|^{3}\cdot\Gamma(V_{k},\boldsymbol{\alpha})\leq(const.)\widetilde{\alpha}^{-1}k^{-1}\Theta_{\max}|V_{k}|^{3}+\left(\Theta_{\max}-\Theta_{\min}\right)|V_{k}|^{3}\ .$ (15)
Now, taking into account that $\Theta_{\max}\sim k^{-2}$ and $\Theta_{\max}-\Theta_{\min}\lesssim k^{-3}$ completes the proof; note here that $\lambda_{0}(V_{m},\infty\delta_{0})=2-2\cos(\pi/|V_{m}|)$, $m\in\mathds{N}$; see, for example, [KY25].
We now turn to (ii): Theorem 5 and Lemma 6 imply for all but finitely many $k\in\mathds{N}$,
$|V_{k}|^{3}\cdot\Gamma(V_{k},\boldsymbol{\alpha})\geq(const.)\widehat{\alpha}^{-1}\lambda_{0}(V_{k},\infty\delta_{0})k^{-1}|V_{k}|^{3}+\left(\lambda_{0}(V_{k},\infty\delta_{0})-\Theta_{\max}\right)|V_{k}|^{3}\ .$ (16)
By direct calculation one has $\Theta_{\max}-\lambda_{0}(V_{k},\infty\delta_{0})\lesssim k^{-3}$. Note that the last term is negative whenever $J\neq\{0\}$ and it vanishes if and only if $J=\{0\}$. Therefore, taking (5) into account, we see that $\widehat{\alpha}$ has to be sufficiently small for $J\neq\{0\}$ in order to obtain a non-trivial lower bound. (5) also implies the statement. ∎
Remark 8.
We now offer a more detailed analysis and interpretation of the upper and lower bounds obtained in Theorem 7 and its proof.
  1. (i)
    Regarding the upper bound (15): Note that the first term is of the form $(const.)\widetilde{\alpha}^{-1}$ and consequently converges to zero when the potential strength converges to infinity in the sense that $\widetilde{\alpha}\to\infty$. However, the last term is independent of the potential strength and is non-zero if and only if $r_{\max}\neq-r_{\max}$. Therefore, the entire upper bound is of the form $(const.)\widetilde{\alpha}^{-1}$ and consequently converges to zero as $\widetilde{\alpha}$ converges to infinity if and only if $r_{\max}=-r_{\min}$.
  2. (ii)
    Regarding the lower bound (16): The last term is negative whenever $J\neq\{0\}$ and it vanishes if and only if $J=\{0\}$; note that, trivially, $\Gamma(V_{k},\boldsymbol{\alpha})\geq 0$ for all $k\in\mathds{N}$. Therefore, this lower bound is non-trivial if and only if $\widehat{\alpha}$ is sufficiently small or if $J\in\{0\}$. Furthermore, the first term of this bound is of the form $(const.)\widehat{\alpha}^{-1}$. Consequently, if and only if $J=\{0\}$, the entire lower bound is of the form $(const.)\widehat{\alpha}^{-1}$ and converges to zero as the strength of the potential tends to infinity in the sense that $\widehat{\alpha}\to\infty$. Note that the lower bound converges to $\infty$ as $\hat{\alpha}\to 0$. This refers to the case where we approach the free Laplacian, for which the spectral gap vanishes in lower order.
  3. (iii)
    The reason we end up with a trivial lower bound when $J\neq\{0\}$ and $\widehat{\alpha}$ is too large is the lack of a sufficiently good lower bound for $\lambda_{1}(V_{k},\boldsymbol{\alpha})$ in terms of $\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0})$ and/or $\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})$; compare with Lemma 6. Indeed, also from the physical point of view outlined in Remark 4, it is reasonable to expect that the lower bound for $\lambda_{1}(V_{k},\boldsymbol{\alpha})$ in Lemma 6 is suboptimal for strong external potentials. On the other hand, however, a corresponding statement as in $(ii)$ of Theorem 7 is expected for all potential strengths.
We conclude this paper with the following corollary which summarizes our findings for the special case $J=\{0\}$ addressed in [KY25], and which originally motivated this work. It also proves the conjecture put forward in [KY25, Section 4].
Corollary 9 (Asymptotic upper and lower bounds II).
Consider the Hamiltonian $H_{k,\alpha}=\mathcal{L}_{k}+\alpha\delta_{0}$ with any $\alpha>0$ and associated spectral gap $\Gamma(V_{k},\alpha)$. Then there exist constants $c_{1},c_{2}>0$ independent of $k$ and $\alpha$ such that for all but finitely many $k\in\mathds{N}$,
$\dfrac{c_{1}}{\alpha}\leq|V_{k}|^{3}\cdot\Gamma(V_{k},\alpha)\leq\dfrac{c_{2}}{\alpha}\ .$
Proof.
At first, with Theorem 2 and (4) we conclude for all but finitely many $k\in\mathds{N}$,
$|V_{k}|^{3}\cdot\Gamma(V_{k},\alpha)\leq(const.)\alpha^{-1}k^{-1}\lambda_{0}(V_{k},\infty\delta_{0})|V_{k}|^{3}\ .$
Secondly, Theorem 5 and again (4) imply
$|V_{k}|^{3}\cdot\Gamma(V_{k},\alpha)\geq(const.)\alpha^{-1}k^{-1}\lambda_{0}(V_{k},\infty\delta_{0})|V_{k}|^{3}$
for all but finitely many $k\in\mathds{N}$. In a final step, recall (5).
Let us remark that, in the continuous one-dimensional setting working with operators on an interval, a result similar to Corollary 9 was established in [KT24] for the special case of a symmetric step potential. To the best of our knowledge, this is so far the only potential for which such bounds have been derived. On the other hand, in [Ker22], the authors could prove corresponding lower bounds for a larger class of symmetric and compactly supported potentials but those bounds are quartic in the length and do not, as conjectured in [KT24], involve the length to the power three.